OpenStudy (anonymous):

Convert the rate 30 cm/s to m/h

OpenStudy (anonymous):

30 cm is 0.3 m 1 s is 1/3600 h hence 30 cm/s = 0.3/(1/3600) m/h = 0.3 * 3600 m/h = 3 * 360 = 1080 m/h

OpenStudy (anonymous):

Okay, thanks.